OpenStudy (anonymous):

7x-29-21x=3-(12+2x) help.

OpenStudy (anonymous):

Finding x?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Okay. 7x-29-21x=3-(12+2x) Combine like terms: -14x-29=3-12-2x -14x-29 = -9-2x -12x-29 = -9 -12x = 20 Divide by -12. x= -5/3

OpenStudy (anonymous):

If you need me to explain more, I will :)

OpenStudy (anonymous):

thank you :)